Determine the result and the propagation of error in the following equation.

(2.8542 + 0.0070) + (2.431 + 0.100)

------------------------------------------

(1.4915 + 0.0069) x (3.459 + 0.049)

Record your answer in the essay (next question) for the error write it as +/-, so your answer should be written like 3.4563 +/0.0034
Answer #1

Let c = \frac{(2.8542 \pm 0.0070) + (2.431 \pm 0.100)}{(1.4915 \pm 0.0069) \times (3.459 \pm 0.049)}

Let n = (2.8542 \pm 0.0070) + (2.431 \pm 0.100) and d =(1.4915 \pm 0.0069) \times (3.459 \pm 0.049)

Then,

n = 2.8542 + 2.431 = 5.2852 and \Delta n = \sqrt{0.0070^2+0.100^2} = 0.1002

d = 1.4915 \times 3.459 = 5.1591 and \Delta d = d\sqrt{(\frac{0.0069}{1.4915})^2+(\frac{0.049}{3.459})^2} = 5.1591\times 0.0149 = 0.0769

c = \frac{5.2852}{5.1591} = 1.0244 and \Delta c = c\sqrt{(\frac{0.1002}{5.2852})^2+(\frac{0.0769}{5.1591})^2} = 1.0244\times 0.0241 = 0.0247

Therefore, result is 1.0244 \pm 0.0247